A solid sphere is rollling without slipping,down,an inclined plane. If the rotational kinetic energy of sphere is 5Joules. Then the total kinetic energy (in Joules) of the sphere is?

Grade:11

1 Answers

Parul singh
7 years ago
Total kinetic energy= 1/2Iw^2 +1/2mv^21/2Iw^2= 1/2*2/5mv^2=5 1/2mv^2=25/2Total kinetic energy=5+25/2=17.5